### 04 Mar 2026 v1

# Impact of 5G SA Logical Vulnerabilities on UAV Communications: Threat Models and Testbed Evaluation

Abstract
This paper studies the impact of logical vulnerabilities in 5G Standalone networks on UAV command and control communication. The work considers three attacker positions in the architecture: a malicious UE connected to the same logical network as the UAV, an adversary with access to the 5G core, and a compromised gNodeB. To evaluate these scenarios, a testbed was built using Open5GS, UERANSIM, and Kubernetes. The environment reproduces a UAV–GCS communication system over a 5G SA network and allows controlled execution of attacks on different network interfaces. The experiments show that attacks at different architectural points can disrupt UAV operations, including manipulation of control commands and termination of data sessions. The results highlight the importance of isolation mechanisms in the 5G user plane and integrity protection in UAV command protocols.
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2027 San Diego, CA, USA https://dx.doi.org/10.14722/ndss.2027.xxxxx www.ndss-symposium.org
I Introduction
Unmanned aerial vehicles (UAVs) rely on continuous communication with the control station to execute commands and transmit telemetry. In scenarios connected by cellular networks, this communication becomes directly dependent on the functioning of the network infrastructure.
5G Standalone networks offer a communication environment with low latency and high transmission capacity. These characteristics allow the transport of control traffic and data between the UAV and the control station using the mobile network infrastructure itself.
Despite these properties, communication between applications and the network involves multiple architectural elements. Connected devices, network core functions, and access nodes participate in the transport and management of data sessions. Each of these points can influence communication behavior.
This work analyzes how attacks conducted at different positions in the 5G architecture can affect UAV command and control communications. Three threat models are considered: a malicious device on the same logical network, an adversary with access to the network core, and a compromised access node.
To evaluate these scenarios, a testbed based on Open5GS, UERANSIM, and Kubernetes was built. The environment replicates a 5G Standalone network and allows observation of the effects of attacks on communication between the UAV and the control station.
II Mobile Networks and 3GPP 5G
The 5G Standalone architecture defines a clear separation between control plane and user plane functions. At the core of the network are functions such as AMF, SMF, UPF, UDM, AUSF, and PCF, organized according to the Service Based Architecture model. Network access is achieved through gNodeB base stations, which connect user devices to the 5G core.
The control plane is responsible for mobility management, authentication, and session establishment. The user plane transports data between devices and the external data network through GTP-U tunnels. The creation and management of these tunnels is performed by the PFCP protocol between the SMF and UPF.
The architecture also introduces concepts such as network slicing and Data Network Name (DNN), which allow the creation of logical communication domains. Devices that establish sessions on the same DNN can share the same data domain depending on the isolation policies configured on the network. In certain scenarios, this configuration can allow direct communication between devices connected to the same slice.
These characteristics influence the system’s attack surface, as different points in the architecture can be exploited to affect communication between devices connected to the network.
III UAV–GCS Scenario in the Testbed
To evaluate attack scenarios, an experimental environment was developed that replicates a standalone 5G network in a controlled environment. The testbed was implemented using a Kind-based Kubernetes cluster, where the 5G core components were deployed in containers.
The network core uses the Open5GS implementation, including the AMF, SMF, and UPF functions. The access layer is simulated with UERANSIM, which provides gNodeB and UE device implementations. The environment uses a single PLMN configured with MCC 999 and MNC 70 and a single slice defined by SST 1 and SD 0x111111.
Three UE devices are used in the experimental scenario. The first represents the UAV, which runs a MAVLink simulator responsible for telemetry and control communication. The second represents the ground control station (GCS). The third device represents a malicious agent used in the experiments.
Depending on the threat model analyzed, the devices may be connected to the same gNodeB or to different gNodeBs. In specific scenarios, an additional pod is also created within the core network with access to SMF and UPF functions to simulate an internal adversary with access to the N4 interface.
This environment allows for the replay of attacks at different points in the 5G architecture and observation of their effects on communication between UAVs and GCS.
IV Threat Models and Attacks
To analyze the impacts of logical attacks on the 5G network upon UAV-GCS communication, we considered three types of scenarios, divided as follows: (1) a malicious UE connected to the same network, targeting the UAV-GCS link; (2) an insider with access to the 5G network core (e.g., N4); and (3) a compromised gNodeB where the attacker fully controls the interface between the UEs and the core. For each of these threat models, the attacker’s capabilities, assumptions, and assets at risk will be described, along with an example attack in our testbed.
IV-A Threat Model 1: Rogue UE Impersonating the UAV
As established by the 3GPP architecture, the notions of slice (S-NSSAI), DNN, and 5G VN Group are distinct concepts that are important for understanding this threat scenario. A slice is a logical partition of network functions and resources. A DNN identifies the data network to which a PDU session is established. In typical implementations, UEs that establish PDU sessions to the same DNN are connected to the same data network and may share the same addressing space, depending on the forwarding and isolation policies configured in the user plane. Thus, the 5G architecture by itself does not guarantee isolation between UEs that establish PDU Sessions to the same DNN. In this context, scenarios can arise in which multiple UEs share the same data domain, leaving it to the system operator to introduce additional configurations, as allowed by the 3GPP specification, for isolation and filtering between UEs or within the data network itself.
For 5G LAN-type services, the 3GPP specification defines a 5G VN Group as a set of UEs that can communicate with each other and provides for unicast traffic forwarding between two UEs within the same 5G VN group. Thus, one UE can send traffic to another UE via the user plane (e.g., local switching at the UPF). Slice security is focused on isolation between slices.
In a context where UEs are in the same slice (intra-slice), user devices can be regarded as accessible attack vectors, as they are outside organizational protection and may be operated by non-technical users. As discussed in, unauthorized access to a slice or its services can create an environment that is vulnerable to various attacks. Moreover, the very attachment to the slice can lead to privacy issues, for example by allowing correlation of users based on slice identification.
Based on these concepts of slicing, DNN, and 5G VN Group, a malicious UE that connects to the same slice as the UAV (and typically to the same DNN) through identity spoofing, impersonation, or failure of UAS-specific authentication or authorization mechanisms, can place itself on the same logical network and, depending on UPF configuration, in the same addressing space. The attacker can therefore send commands directly to the UAV.
In this context, an attacker might consider using lateral movement techniques and traditional IP-based exploits, such as network scanning and direct exploitation of the target UE’s (UAV) IP address.
With this threat model in mind, a script was created to start the testbed with the following configuration: (1) the PLMN is configured for MCC 999 and MNC 70; (2) a single slice is used: SST 1 and SD 0x111111, defined in the AMF and subscribers; (3) the DNN is of type IPv4 with APN ”internet” where all three UE’s share the same logical network (VN Group / slice / DNN); and (4) the UAV, GCS, and ROGUE are attached to the same gNodeB.
Figure 1 shows the architecture of the testbed with this threat model:
Figure 1: Threat Model 1: UAV, GCS, and ROGUE UE attached to the same gNodeB.
To explore this scenario and demonstrate the impact on UAV operations, we implemented a C2 Injection attack at the application layer, also taking advantage of the fact that the main UAV applications that use MAVLink, such as ArduPilot, PX4, and QGroundControl, for example, do not use signing as a standard in their applications.
Since in this threat model the ROGUE UE can directly reach the legitimate UAV, because it is in the same slice and DNN, the attacker performs a network scan to identify the UAV (e.g., by probing port 14550), as seen in Listing  1. From this and targeting these vulnerabilities, the ROGUE UE can impersonate the control station by configuring the MAVLink system identifier as 255 (Ghost GCS), the default value associated with the GCS.
Listing 1: UDP scan for MAVLink port 14550 from ROGUE UE pod.
root@ueransim-rogue-6677bfc74d-7s2pm:/tmp# nmap -sU -p 14550 -Pn -n 10.45.0.2
Starting Nmap 7.95 at 2026-03-03 18:46 UTC
Nmap scan report for 10.45.0.2
Host is up.
PORT STATE SERVICE
14550/udp open|filtered unknown
Nmap done: 1 IP address (1 host up) scanned in 2.98 seconds
The attack is divided into two phases: (1) soliciting UAV telemetry and obtaining its position (2) flooding the UAV with commands.
In the first phase, considered reconnaissance and espionage, ROGUE attempts to establish a ”MAVLink” connection on UDP port 14550 towards the UAV, sending heartbeats impersonating the legitimate GCS and using the command MAV_CMD_SET_MESSAGE_INTERVAL to request the UAV to send the message GLOBAL_POSITION_INT, which returns its position and is necessary for the flooding algorithm to function. If successful, the attacker can obtain its real position (latitude and longitude) as well as other forms of telemetry, compromising the confidentiality of the UAV from the outset. Even if there is no response (for example, firewall or application restrictions), the algorithm still proceeds to the flooding phase.
In the second phase, the algorithm sends a flood of MAV_CMD_NAV_LAND commands to force the UAV to land, creating a race condition with the legitimate GCS. Due to the high rate of command transmission, the chance of the UAV accepting malicious commands increases, resulting in the forced landing of the aircraft and the neutralization of the mission, thus compromising the three pillars of information security: confidentiality, integrity, and availability.
In the testbed, to simulate a real attack, the attacker directly accesses the Kubernetes pod of the ROGUE UE and executes the ‘nmap‘ application to scan the network and then execute the exploit on the identified port/IP. The script, as per Algorithm 1, sends the heartbeat and the telemetry request. The algorithm then waits a few seconds and starts sending the MAV_CMD_NAV_LAND commands.
The result observed in our tests in the testbed is that the UAV accepted the commands, forcing its eventual landing and end of the mission, regardless of the legitimate operator’s intentions(Listings 3). The failsafe based on the GCS heartbeat timeout does not activate, as the ROGUE also sends heartbeats as GCS. The central problem is the lack of authentication in the application, not the loss of the link.
Algorithm 1 C2 Injection and GCS Spoofing
1:procedure C2-INJECTION-GCS-SPOOFING(targetIP, targetPort)
2:  $master\leftarrow\text{MAVLinkConnection}(\text{udpout}:\text{targetIP}:\text{targetPort},\text{source\_system}\leftarrow 255)$ $\triangleright$ Impersonate GCS (system ID 255)
3:  Send(mastermaster, HEARTBEAT(MAV_TYPE_GCS, MAV_AUTOPILOT_INVALID)) $\triangleright$ Establish “session” toward UAV
4:  $lat\leftarrow 0$; $lon\leftarrow 0$; $mode\leftarrow\text{``PRECISE''}$ $\triangleright$ Initialize attack state
5:  Send(mastermaster, COMMAND_LONG(SET_MESSAGE_INTERVAL))
6:  $startTime\leftarrow\text{now}()$; $timeout\leftarrow 5$
7:  while (now()startTime)<timeout(now()-startTime)<timeout do
8:   $msg\leftarrow\textsc{RecvMatch}(master,\text{type}\leftarrow\text{GLOBAL\_POSITION\_INT})$
9:   if $msg\neq\emptyset$ then
10:     $lat\leftarrow msg.\text{lat}/10^{7}$; $lon\leftarrow msg.\text{lon}/10^{7}$; $alt\leftarrow msg.\text{relative\_alt}/1000$ $\triangleright$ Exfiltrate UAV position (confidentiality breach)
11:     break
12:   end if
13:  end while
14:  if $lat=0\wedge lon=0$ then
15:   $mode\leftarrow\text{``BLIND''}$ $\triangleright$ No position received; force vertical land in place
16:   for $i\leftarrow 1\textbf{ to }50$ do
17:     Send(mastermaster, COMMAND_LONG(MAV_CMD_NAV_LAND, $0,0,0,0,0,lat,lon,0$)) $\triangleright$ Phase 2: flood LAND commands (integrity/availability breach)
18:     Sleep(0.04) $\triangleright$ Accelerated rate to win race against legitimate GCS
19:   end for
20:  end if
21:  return
22:end procedure
Listing 2: Execution of the C2 injection exploit: BLIND attack with MAV_CMD_NAV_LAND commands
root@ueransim-rogue-66f7bfc74d-7s2pm:/tmp# /opt/rogue-venv/bin/python3 exploit.py 10.45.0.2
[*] TARGET SET: 10.45.0.2:14550
[*] Connecting as Secondary GCS (ID 255)...
[*] Attempting to request GPS stream via MAV_CMD_SET_MESSAGE_INTERVAL...
[.] Waiting for telemetry data for 5 seconds...
[-] Drone ignored the data request (Possible UDP restriction).
[!] SWITCHING STRATEGY TO: BLIND ATTACK.
[!] The drone will be forced to land at current vertical position.
[*] Starting neutralization sequence (Mode: BLIND)...
[>] Firing command packet 0/50...
[>] Firing command packet 10/50...
[>] Firing command packet 20/50...
[>] Firing command packet 30/50...
[>] Firing command packet 40/50...
[SUCCESS] Commands sent. Check UAV status (Landing Initiated).
Listing 3: Threat Model 1: logs of the UAV state after the attack (forced landing).
... (repeated)
[UAV] heartbeat sent
[UAV] heartbeat sent
[UAV] heartbeat sent
[UAV] landed.
[UAV] heartbeat sent
[UAV] heartbeat sent
[UAV] heartbeat sent
... (repeated)
To mitigate this vulnerability, we recommended to activate and use the signing function provided by MAVLink for the application layer. This will render the ROGUE impersonation attack using code 255 insufficient to allow unauthorized commands to be accepted.
IV-B Threat Model 2: Insider Threat
This threat model considers an insider adversary without radio equipment or UE registration, but with access to the 5G core network. In this scenario, it is assumed that the attacker reached the core (e.g., via compromised infrastructure, supply chain, or misconfigured isolation) and has access to both SMF and UPF, in particular, access to the N4 interface (PFCP, UDP port 8805), an environment already depicted in [1]. On this interface, SMF sends commands to UPF to create, modify, and delete PDU Sessions and the associated GTP-U tunnels [8]. In this context, messages can be manipulated to disrupt UE connectivity on the User Plane [1][6].
In practice, N4 and other internal non-SBA interfaces may not be cryptographically protected: 3GPP allows the operator to rely on other means (e.g. physical or logical trust) and leaves it to operator choice whether to use IPsec or other cryptographic protection on interfaces such as N4 and N9 [9]; when interfaces are considered trusted (e.g. physically protected), the operator may decide not to use cryptographic protection [9]. Prior work has demonstrated attacks under sub-optimal or inadequate security of the N4 interface [1], and N4 is often described as a trusted interface in 5G security analyses [7], which aligns with deployments where N4 is not fully secured.
Figure 2: Threat Model 2: Insider Adversary with Access to the 5G Core Network.
By gaining access to the N4 interface, the attacker can execute several attacks: (i) spoofing an SMF record from the UPF and, if successful, sending control messages to the UPF to disrupt the connection of one or more UEs (e.g., Session Deletion Request, Session Modification Request); (ii) replaying captured PFCP traffic; and (iii) flooding messages such as session establishment, modification, or deletion requests. Furthermore, it is assumed that the attacker does not need to break any cryptographic protection or possess subscriber credentials. The threat is confined to the core with the premise that the N4 interface is not protected and it is possible to inject commands into the UPF [1]. This setup reflects the reality that core networks are often assumed to be trusted and may lack cryptographic protection or strict segmentation on internal interfaces [9][7], making the presence of an attacker with access to the core a plausible scenario.
To emulate this model, an initialization script was created with the parameters shown above and with the following configuration: (1) the PLMN is configured for MCC 999 and MNC 70; (2) a single slice is used: SST 1 and SD 0x111111, defined in the AMF and subscribers; (3) the DNN is of type IPv4 with APN ”internet” where all three UEs share the same logical network (VN Group / slice / DNN); and (4) the UAV, GCS, and ROGUE are attached to the same gNodeB. In addition, a pod was created in Kubernetes within the core’s logical network, with access to the SMF and UPF functions (N4 interface), as shown in Figure 2.
With the aim of demonstrating a possible impact on drone operations using this threat model, we implemented an algorithm that generates a flood of messages to delete (Session Deletion Request), incrementing the SEID with each loop, thus creating a black-box attack on the UPF and applying and improving procedures already described in other papers on PFCP [1].
In our experiment, instead of launching the attack directly from the SMF, we opted for a more realistic approach, where the attacker (insider) executes the script that first performs a PFCP Association with the UPF, thereby registering as an additional control-plane peer alongside the legitimate SMF.
After the association is established with the UPF, the script initiates the following phases: (i) iteration over a range of SEIDs passed as parameters (brute-force enumeration); (ii) for each SEID, a Session_Deletion_Request with a unique sqn is sent and the deletion response is analyzed; and (iii) when the response indicates Cause = Request Accepted, the corresponding PDU Session is terminated.
The result observed in the experiments is the teardown of the PDU Session (and its GTP-U tunnel) for each SEID for which the UPF returned Request Accepted. Regardless of whether only the UAV or GCS session was interrupted, the observed behavior was the UAV entering failsafe mode and returning to its point of origin, impacting its operations.
Mitigations for this threat include protecting the N4 interface with mutual authentication and integrity (e.g., IPsec or TLS where supported), so that the UPF accepts PFCP messages only from authenticated peers [9]. Operators can restrict which Node IDs are allowed to establish PFCP associations and send session-level commands, and segment the core network so that only designated hosts can reach the SMF and UPF [1]. Monitoring and anomaly detection on N4 (e.g., bursts of Session Deletion Requests or traffic from unexpected sources) can help identify insider or compromised-node abuse [7].
Algorithm 2 PFCP Session Deletion Flood
1:procedure PFCP-SESSION-DELETION-FLOOD(SMFaddress, UPFaddress)
2:  $\triangleright$ Execution from SMF or from insider after Association Setup
3:  $SEID\leftarrow 1$ (or seid_startseid\_start); $delCounter\leftarrow 0$
4:  while $SEID\leq seid\_max$ (or for each SEIDSEID in seid_listseid\_list) do
5:   $sqn\leftarrow\text{next unique sequence number}$
6:   $pktPayload\leftarrow\textsc{SessionDeletionRequest}(SEID,sqn)$
7:   SendRequest($src\leftarrow SMFaddress$, $dst\leftarrow UPFaddress$, pktPayloadpktPayload)
8:   (optionally) wait for SessionDeletionResponse
9:   if $\text{Cause}=\text{Request Accepted}$ then
10:     $delCounter\leftarrow delCounter+1$
11:     $\triangleright$ PDU session torn down; UAV–GCS tunnel may be lost
12:   end if
13:   $SEID\leftarrow SEID+1$ (or next in list)
14:  end while
15:  return delCounterdelCounter
16:end procedure
IV-C Threat Model 3: Compromised gNodeB
This threat model considers an adversary that has gained control of a gNodeB (gNB) that is already integrated into the legitimate 5G Standalone network. Unlike an external fake base station, the compromised gNB remains part of the architecture: it participates in the Service-Based Architecture (SBA) and in the user plane (GTP-U), so it can observe and manipulate both control-plane signalling (N2/NAS) and the Command and Control (C2) traffic of UEs attached to it. Compromise may occur through realistic vectors such as exploitation of known vulnerabilities (CVEs) in virtualized RAN or O-RAN components, unauthorized access to management interfaces (e.g., SSH, NETCONF, O1/OAM), supply-chain attacks in cloud-native deployments, or failures of isolation between network functions in shared infrastructure [4]. In private 5G networks used for critical operations (e.g., BVLOS UAV missions), the larger attack surface of the access node makes a logically compromised gNB a plausible and relevant internal threat [7].
Why the gNB is a critical point. The gNB is the termination point of the radio link ciphering (PDCP). Even when the air interface is protected (e.g., NEA2), the gNB must decapsulate and decrypt user-plane packets to forward them toward the core. Thus, inside a legitimate gNB, user data (including the inner IP and application payload) is available in cleartext. The 3GPP specifications do not mandate encryption or integrity protection on the N3 interface (GTP-U between gNB and UPF): protection is often left to operator choice, and in practice many deployments do not use IPsec or integrity on N3 [9]. Operators are reluctant to deploy IPsec on the N3 interface because it is CPU-intensive and reduces the throughput of user traffic; moreover, user data is often perceived to be protected at the application layer (e.g., TLS), leading some to consider IPsec redundant [2]. While 3GPP recommends securing the N2 (NGAP) and N3 (GTP-U) interfaces using IPSec, encryption is not mandatory, only integrity protection is strongly advised, leaving these interfaces accessible for monitoring in many deployments [5]. Studies report that GTP-capable nodes, including UPFs, are in some cases exposed to the public Internet, contradicting the ”walled garden” assumption: measurements have identified hundreds of thousands of valid GTP hosts accessible via the public Internet across multiple countries and providers [11]. For the user plane on the radio and backhaul segments, integrity protection is frequently optional (e.g., for fronthaul and N3): integrity protection of fronthaul traffic is optional in the protocol standards due to performance requirements of the eCPRI protocol, and lack of integrity over fronthaul is deemed acceptable by some specifications under perceived low risk [10]; empirical studies of live deployments show that User Plane integrity (e.g., NIA2) is often absent, with only a minority of operators enabling it [3, 4]. Together, this increases the risk that a compromised gNB can read and modify traffic without the network or the UE detecting it. In summary: (1) the gNB terminates radio encryption and thus has cleartext access to the payload; (2) the N3/GTP-U path often has no cryptographic integrity; (3) a compromised gNB can therefore tamper with user-plane traffic, including C2 protocols carried inside the tunnel.
Attack: UAV Navigation Hijacking via Compromised gNB Data Tampering (ATK-ROGUE_GNB-01). We consider the case where the UAV is served by a legitimate gNB (gNB01) and the GCS is served by a compromised gNB (gNB-rogue). All C2 traffic from the GCS to the UAV traverses the 5G core and, on the return path to the UAV, may pass through the compromised gNB if the topology so dictates; in our testbed we arrange so that the GCS is attached to the rogue gNB, so that uplink C2 (GCS → UAV) crosses the compromised node. The attacker’s goal is to alter navigation commands without breaking the tunnel or the radio link. MAVLink messages such as SET_POSITION_TARGET_LOCAL_NED carry waypoint or position targets; they are transported in cleartext inside the GTP-U tunnel. A compromised gNB that can inspect and modify GTP-U payloads can replace the target coordinates with values chosen by the attacker, causing the UAV to fly toward an unintended location while the operator believes the original command was executed. This corresponds to MITRE ATT&CK Impact (TA0040) and the technique T1565 (Data Manipulation).
To emulate this model we use a custom script, which deploys a 5G SA environment in a Kubernetes (Kind) cluster with Open5GS (AMF, SMF, UPF, etc.) and security hardening on the AMF (NIA0/NEA0 removed). The user plane is left without N3 encryption/integrity, consistent with 3GPP optionality and common practice. There are three UERANSIM gNBs: gNB01 (nci 0x10), gNB02 (nci 0x20, no UEs), and gNB-rogue (nci 0x30), the latter under attacker control but still connected to the legitimate core. The UAV (IMSI 001) attaches only to gNB01 and runs a MAVLink simulator; the GCS (IMSI 002) and a ROGUE UE (IMSI 003) attach to gNB-rogue. Thus, all GCS traffic, including C2 toward the UAV, passes through the compromised gNB. PLMN is 999/70 and a single slice (SST 1, SD 0x111111) is used.
[…]
Observed result. In the testbed, when the GCS sends a position target, the modified command causes the UAV to move toward the attacker’s chosen coordinates. The operator sees the command accepted by the UAV but not the altered target, demonstrating navigation hijacking via data manipulation. The vulnerability is cross-layer: the 5G user plane often does not provide integrity for N3/GTP-U, and the C2 application (MAVLink) relies on that channel without end-to-end integrity (e.g., signing). A compromised gNB exploits both aspects.
Mitigation. Mitigations align with the evidence from the literature and slides: (i) User-plane integrity: where feasible, enable integrity (and optionally confidentiality) on the N3 segment (e.g., IPsec between gNB and UPF) or future 3GPP mechanisms for user-plane integrity, so that tampering by a compromised node becomes detectable. (ii) Application-layer protection: use MAVLink signing (or equivalent integrity/authenticity) so that the UAV verifies origin and integrity of C2 commands before execution; this protects even if the 5G path is compromised. (iii) Monitoring and segmentation: monitor C2 traffic for anomalies at gNB/UPF and isolate C2 in a dedicated slice or session with stricter policies, reducing the impact of a single compromised gNB.
V Conclusion
This work analyzed the impact of logical vulnerabilities in 5G Standalone networks on communication systems between UAVs and control stations. Three threat models were considered, representing distinct attacker positions within the network architecture.
Experiments conducted in the testbed demonstrated that attacks carried out by a malicious device connected to the same slice, by an adversary with access to the network core, or by a compromised gNodeB can directly affect UAV operation. Depending on the scenario, interruption of data sessions, manipulation of navigation commands, and execution of malicious commands were observed.
The results indicate that the security of UAV communication depends on mechanisms distributed across different layers of the system. Network configurations that allow direct communication between devices in the same data domain can facilitate lateral attacks. Internal core interfaces, when not protected by cryptographic mechanisms, can allow manipulation of sessions at the user plane. Application protocols that do not use authentication or integrity also amplify the impact of these attacks.
These observations reinforce the need to consider multi-layered protection in the architecture, including isolation mechanisms in the 5G network and integrity protection in the UAV control protocol.